OpenStudy (anonymous):

Simplify: 3^3 / 9 x 2 - 12 + 6

OpenStudy (amistre64):

9/9=1*2 = 2-12 = -10+6 = -4 ??

OpenStudy (anonymous):

Wow, you're quick..

OpenStudy (amistre64):

quick yeah, but blind

OpenStudy (amistre64):

27/9=3*2=6-12=-6+6=0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Wow, you guys are amazing!! :)

OpenStudy (amistre64):

.... and i can knit :)

OpenStudy (anonymous):

I can't...my mother didn't teach me :(

OpenStudy (anonymous):

i have given the equation

OpenStudy (anonymous):

I apparently can't knit OR do math! :D Thanks everyone!

OpenStudy (amistre64):

some of us, i wont name names, are over acheivers ;)

OpenStudy (anonymous):

Nothing wrong with that! ;-)